OpenStudy (anonymous):

How far have you traveled if you drive 5 1/2 hours at 59 mph? Round to the nearest mile. A. 300 miles B. 325 miles C. 350 miles D. 375 miles

OpenStudy (anonymous):

The distance you travel is equal to your velocity times time: \[d=vt=59 mph \times 5.5 hours = 324.5 miles \approx 325 miles\]

OpenStudy (mathstudent55):

average speed = distance / time distance = (average speed) * time

OpenStudy (anonymous):

The algebraic methods above are both correct. Also consider the approximation method. \[5.5 hours*60-5.5=330-5.5=324.5 \approx 325\]This method lets you get the answer in your head, without pencil/paper/calculator. Mental math is useful out there in real life....